# BlockLLM: Memory-Efficient Adaptation of LLMs by Selecting and Optimizing the Right Coordinate Blocks

### v1

@@ Abstract
Training large language models (LLMs) for pretraining or adapting to new tasks and domains has become increasingly critical as their applications expand. However, as the model and the data sizes grow, the training process presents significant memory challenges, often requiring a prohibitive amount of GPU memory that may not be readily available. Existing methods such as low-rank adaptation (LoRA) add trainable low-rank matrix factorizations, altering the training dynamics and limiting the model’s parameter search to a low-rank subspace. GaLore, a more recent method, employs Gradient Low-Rank Projection to reduce the memory footprint, in the full parameter training setting. However GaLore can only be applied to a subset of the LLM layers that satisfy the “reversibility” property, thus limiting their applicability. In response to these challenges, we introduce BlockLLM, an approach inspired by block coordinate descent. Our method carefully selects and updates a very small subset of the trainable parameters without altering any part of its architecture and training procedure. BlockLLM achieves state-of-the-art performance in both finetuning and pretraining tasks, while reducing the memory footprint of the underlying optimization process. Our experiments demonstrate that fine-tuning with only less than 5% of the parameters, BlockLLM achieves state-of-the-art perplexity scores on the GLUE benchmarks. On Llama model pretrained on C4444 dataset, BlockLLM is able to train with significantly less memory than the state-of-the-art, while still maintaining competitive performance.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Recent advancements in natural language processing (NLP) have been propelled by the development of large language models (LLMs) [15, 30, 22, 1]. These models have set new benchmarks for a variety of NLP tasks, including language translation [29], text summarization [13], and sentiment analysis [3]. The core strength of LLMs lies in their scale. Empirical evidence suggests that increases in model size not only enhance performance across standard benchmarks but also unlock new capabilities that are absent in smaller models [12, 35]. Pre-training and finetuning LLMs on domain-specific application data have enhanced their applicability immensely.
@@
Existing strategies for memory-efficient training.
@@
To address these challenges, multiple strategies are being explored to reduce the number of parameters, gradients, and the corresponding optimizer state size. One popular strategy is the application of pruning methods, where a large set of parameters or even entire layers are removed from the model architecture [32, 18, 28]. However, pruning approaches often require extensive retraining to recover lost accuracy [5]. Furthermore, identifying which parameters are crucial before training is challenging [20, 26]. This challenge complicates implementation and can lead to generalization issues, particularly on diverse or unseen data [18].
@@
PEFT (Parameter-Efficient-Fine-Tuning) methods such as [9, 16, 10] achieve memory efficiency by introducing low-rank matrices to the transformer architecture, significantly reducing the number of trainable parameters needed during fine-tuning. Although integrating these low-rank matrices alleviates the extensive retraining demanded by pruning techniques, they can alter the training dynamics, potentially leading to quality issues during the merging phase [8]. Moreover, the additional parameters introduced by PEFT methods can increase the model’s parameter size, countering efforts to reduce overall model size.
@@
A recent work, GaLore [34], which focuses on full parameter training, achieves memory efficiency by performing low-rank factorization of the gradients in specific layers. However, GaLore does not achieve high memory efficiency across all model types, as its gradient factorization method can only be applied to layers that satisfy the reversibility property. This limitation restricts its applicability and efficiency in models where not all layers exhibit this property.
@@
Techniques such as gradient and activation checkpointing, quantization, and parameter offloading are also commonly used to achieve memory savings. These methods often come with trade-offs, such as increased computational overhead or compromised performance. For instance, checkpointing [4] reduces memory usage but requires recomputation, quantization lowers precision and can affect accuracy [7] and parameter offloading increases data transfer latency [25].
Block coordinate descent (BCD).
BCD is a popular algorithm in the large-scale optimization literature [21]. At any training iteration $t$, instead of updating all the parameters $W$, BCD selectes only a block of parameters $b_{t}$ and updates only
$W^{b_{t}}_{t+1}=W^{b_{t}}_{t}+d_{t},$
where $d_{t}$ is the update. As it can be seen, this formulation directly falls in the reduced parameter training regime. This insight forms the cornerstorne of our approach and hence the name BlockLLM. A related study by Belilovsky et al. [2] demonstrated that sequentially solving one-hidden-layer problems could match the performance of large model training, inspiring us to tackle parameter and memory efficiency by training large models in parts. By definition, this method doesn’t alter the model architecture in anyway. Furthermore, works such as [24, 33, 21] prove the convergence of BCD on various problem architectures; inspiring us to adapt it to large scale LLM training.
@@
We propose a novel parameter and memory-efficient algorithm, BlockLLM where we dynamically select and train a block of parameters. BlockLLM primarily achieves memory savings by persisting optimizer states only for these selected parameters. We introduce a novel block selection criterion, for LLM training. We demonstrate the memory-efficient training and superior performance of BlockLLM on pre-training and finetuning tasks.
@@ 2 Methodology
They key idea behind BlockLLM is to select and update only a subset of parameters during training, enabling us to achieve significant memory savings. However, the criteria for selecting the “right” subset of parameters is not clear. In this section, we look at magnitude pruning as a tool to identify parameter importance, in the finetuning setting.
@@
Magnitude pruning is a widely recognized technique for reducing the parameter count in neural network models [6]. In this analysis, we use parameter magnitude as a measure of parameter importance and study the impact of training on the selected parameters. First, we trained DistilBERT [27] on the IMDb dataset [19] for sequence classification, achieving an accuracy of 92.02%. We then conducted inference on the GLUE-CoLA dataset [31] without fine-tuning, resulting in a significant drop in accuracy to $47.74\%$. This drop in performance, possibly due to domain shift, encouraged us to use this setup for our analysis.
Next, we performed magnitude pruning on the IMDb pre-trained model at various sparsity levels and fine-tuned these pruned models on the GLUE-CoLA dataset. Let $s$ denote the sparsity level, $W^{t}$ represent the model parameters at iteration $t$, and $n$ be the total number of parameters. For each parameter $w_{i}$ where $i=1,\ldots,n$, we compute $|w_{i}|$. During training, we update only $S=\text{Top}_{k}|W^{0}|$, where $k=n\times(1-s)$. The results of these experiments, detailing the relationship between sparsity and accuracy, are summarized in Table 5.
@@
To further understand the effects of this parameter selection strategy, we analyzed $|W|$, before and after training to understand which weights are updated more frequently and their impact on model performance. Specifically, we compared the initial weights $|W^{0}|$ and final weights $|W^{t}|$ of the model after fine-tuning on the CoLA dataset [31]. Our findings are presented in Figure 1. The histogram on the left of Figure 1 shows $|w_{i}^{t}|$ for all $i$ where $\delta_{i}>\eta$. Here, $\delta_{i}=|w_{i}^{0}-w_{i}^{t}|$ and $\eta$ is the threshold. The histogram on the right of Figure 1 plots the frequency of $\delta$ in the updates, revealing that a large percentage of the updates were minor.
@@
(a) Histogram of the changed parameters
(b) Histogram of $\delta$
Figure 1: Analysis of weight magnitudes of DistilBERT model pretrained on IMDb dataset and finetuned on CoLA dataset.
@@ 2.1 Analysis of Reduced Parameter Training
We conducted this experiment using the same setup: fine-tuning the DistilBERT model on the GLUE-CoLA dataset after pre-training on the IMDb dataset. The results of this experiment are summarized in the Table 1. Additionally, we performed similar experiments on other GLUE [31] datasets, with the results presented in Section A.3
$1-s$ $q$ $m$ Accuracy Matthews Correlation
$0.1$ $0.58$ $1000$ $82.55$ $0.5882$
$0.02$ $0.36$ $1000$ $82.45$ $0.5711$
$0.02$ $0.14$ $4000$ $82.45$ $0.5794$
$0.02$ $0.10$ $6000$ $82.07$ $0.5679$
Table 1: Impact of Update Frequency and Sparsity on CoLA Dataset
Our observations indicate that $s$ decreases, meaning more parameters are updated per iteration, the number of unique parameters $q$ also increases, which results in better model performance. The increasing number of unique updates indicates that different parameters may become important at different stages of training, necessitating a more dynamic approach to parameter updates.
@@
  • Parameter Selection Criteria: Our analysis in Figure 1 indicates that the model frequently updates parameters with lower weight magnitudes. This contradicts the very premise of magnitude pruning. Therefore its not clear if adopting magnitude as a parameter importance criteria will help, in general. In our work, we bank on evidence from prior work on greedy parameter selection strategy [24, 21] and use gradient as the parameter importance criteria.
  • Parameter Selection Frequency: As discussed earlier, determining when to change the set of parameters to update for a given training phase is crucial. In BlockLLM, we developed a strategy that uses the current loss objective value as a signal to decide if the parameter selection needs to be updated.
2.2 BlockLLM:
In this section we introduce BlockLLM, a parameter and memory efficient training method designed to reduce the number of trainable parameters in large language models (LLMs) without compromising training performance. Akin to other parameter-efficient fine-tuning (PEFT) methods such as LoRA [9] and ReLoRA [16], BlockLLM updates only $k$ parameters, at any iteration $t$, where $k\ll n$, $n$ is the total number of parameters. However, the main difference is that BlockLLM optimizes parameter selection by focusing on the most impactful parameters at different stages of the training process. The overall algorithm of BlockLLM is given in Algorithm 1.
@@ 2.2 BlockLLM:
In the context of LLMs, at iteration $t$, the update to the parameter is the processed gradient $\tilde{G}_{t}$, calculated using optimizers such as Adam [14]. Specifically, for any layer $l$ of the model, the update is given by,
$\tilde{G}^{l}_{t}$ $=M^{l}_{t}/\sqrt{V^{l}_{t}+\epsilon}\hskip 8.5359pt\text{where,}\ M^{l}_{t}=\beta_{1}M^{l}_{t-1}+(1-\beta_{1})G_{t}^{l}\hskip 5.69046pt\text{and}\hskip 5.69046ptV^{l}_{t}=\beta_{2}V^{l}_{t-1}+(1-\beta_{2}){G_{t}^{l}}^{2}.$ (1)
@@
Here, $\beta_{1}$ and $\beta_{2}$ are hyperparameters of the optimizer and $G_{t}^{l}$ is the gradient at layer $l$ and $M_{t}$ and $V_{t}$ are the Bias-corrected first moment estimate and Bias-corrected second moment estimate respectievly. During the backward pass, BlockLLM selects the layers with the large $|\tilde{G}_{t}^{l}|$ and updates only those layers, denoted as set $S$. Note that by selecting full layers, we may not achieve the desired sparsity level $s$. Therefore, subsequently, for each selected layer we construct a binary mask to retain only the top $k$ parameters by gradient magnitude:
$\text{mask}[i,j]=\begin{cases}1&\text{if }\|\tilde{G}_{t}^{l}[i,j]\|\geq\tau\\ 0&\text{otherwise},\end{cases}$
where $\tau$ is an estimated threshold, computed by looking at the gradient values of each layer. Then, in every iteration $t$, the selected parameters in layers $l\in S$ are updated using the computed masks as:
$W_{t+1}^{l}=W_{t}^{l}-\eta\left(\text{mask}\odot\tilde{G}_{t}^{l}\right).$ (2)
@@
An illustration of the proposed parameter selection procedure is given in Figure 3. However, there is one caveat with this approach. In the initial training iterations, the gradient estimates are known to be noisy. Additionally, there are several cases such as pre-training and finetuning, with heavy domain shift, where there is very little useful inductive bias. Therefore, using gradients to select important parameters may prove to be detrimental to our cause in the initial few iterations.
@@
To address this challenge, our experiments incorporate layer visit frequency $f$ into the selection criteria. Specifically, for any layer $l\in L$, $f_{l}$ represents the z-score normalized number of times the layer has been selected. Consequently, the layer selection criterion is modified to $|\tilde{G}_{l}|/f_{l}$ which tends to select layers with high gradient norms while also considering layers that have been less frequently chosen in previous iterations. Our experimental results demonstrate that this refined selection criterion enhances performance.
Mask and Memory.
In the current implementation of BlockLLM, masks are binary matrices matching the dimensions of the gradients in their respective layers. Consequently, the memory footprint scales with the number of selected layers, as each layer requires a new mask. Therefore, for our experiments, we consider an approximation where we skip the mask computation step and update all parameters in $S$. The update step is as follows:
$W_{t+1}=\begin{cases}W_{t}^{l}-\eta\tilde{G}_{t}^{l}&\text{if }l\in S,\\ \text{Freeze }W^{l}&\text{otherwise}.\end{cases}$ (3)
Our experimental results demonstrate that this approximate block selection strategy is effective in practice. Nonetheless, we believe that a more refined implementation, capable of achieving a more granular parameter selection, is feasible. We reserve this as a potential area for future exploration.
@@
The natural next question is how many iterations to perform the update (3) with the same set $S$. BlockLLM addresses this by using the loss $\phi$ as a critical signal for determining when to change the parameter selection. Specifically, BlockLLM introduces a hyperparameter, patience $m$. At any iteration $t$, if $\phi_{t}$ exceeds the moving average of losses over the last $m$ iterations, the set $S$ is revised. The detailed parameter selection frequency algorithm is provided in in Algorithm 2.
@@
Next, we explore the memory benefits of training with BlockLLM, which stem primarily from its parameter-efficient training approach. In practice, updating fewer parameters directly reduces the number of gradients and optimizer states that need to be stored in VRAM. With $s\%$ sparsity, BlockLLM should directly reduce the optimizer states by $s\%$ compared to full parameter training. Our empirical analysis corroborates this.
@@
10:         $H=[]$triangleright Reset loss history
@@
12:     // Update the selected params
@@
15:         // Adam update as in (1)
16:         $M_{t}^{l},V_{t}^{l},\tilde{G}_{t}^{l}$ = Adam($M_{t-1}^{l},V_{t-1}^{l},G_{t}^{l}$)
17:         $\tilde{G}^{l}_{t}=\text{mask}\odot\tilde{G}_{t}^{l}$
18:         $W_{t+1}^{l}=W_{t}^{l}-\eta\tilde{G}_{t}^{l}$
@@
11:// Compute mask
12:     for each $l\in S$ do
13:         Let $\text{mask}_{l}$ $=\mathbf{0}_{|G_{t}^{l}|}$
14:         for each $i,j$ in mask do
15:              if $\|\tilde{G}_{t}^{l}[i,j]\|\geq\tau$ then
16:                  $\text{mask}_{l}[i,j]=1$
17:              end if
@@
19:     end for
20:     return mask, $S$
21:end function
@@
Figure 3: Given a large language model consisting of many layers, BlockLLM first finds the layers with the largest gradient $||\tilde{G}_{l}||$ (highlighted in orage) and selects a subset of the parameters. During optimization, only the selected layers will be updated ($\theta$) and the Adam optimizer will only keep track of the bias corrected moments ($M_{l}$) and ($V_{l}$) for those selected parameters (shown in orange).

### v2

@@ Abstract
Training large language models (LLMs) for pretraining or adapting to new tasks and domains has become increasingly critical as their applications expand. However, as model and data sizes grow, the training process presents significant memory challenges, often requiring a prohibitive amount of GPU memory that may not be readily available. Existing methods such as low-rank adaptation (LoRA) add trainable low-rank matrix factorizations, altering the training dynamics and limiting the model’s parameter search to a low-rank subspace. GaLore, a more recent method, employs Gradient Low-Rank Projection to reduce the memory footprint, in the full parameter training setting. However GaLore can only be applied to a subset of the LLM layers that satisfy the “reversibility” property, thus limiting their applicability. In response to these challenges, we introduce BlockLLM, an approach inspired by block coordinate descent. Our method carefully selects and updates a very small subset of the trainable parameters without altering any part of its architecture and training procedure. BlockLLM achieves state-of-the-art performance in both finetuning and pretraining tasks, while reducing the memory footprint of the underlying optimization process. Our experiments demonstrate that BlockLLM achieves superior performance on finetuning both large and small models. On pretraining a Llama model on C4444 dataset, BlockLLM is able to train with significantly less memory than the state-of-the-art, while still maintaining competitive performance.
@@ 1 Introduction
Recent advancements in natural language processing (NLP) have been propelled by the development of large language models (LLMs) Le Scao et al. (2023); Touvron et al. (2023); OpenAI (2023); Almazrouei et al. (2023). These models have set new benchmarks for a variety of NLP tasks, including language translation Takase & Kiyono (2021), text summarization Kedia et al. (2021), and sentiment analysis (Brown et al., 2020). The core strength of LLMs lies in their scale. Empirical evidence suggests that increases in model size not only enhance performance across standard benchmarks but also unlock new capabilities that are absent in smaller models (Kaplan et al., 2020; Zhao et al., 2023). Pretraining and finetuning LLMs on domain-specific application data have enhanced their applicability immensely.
@@
Figure 1: Illustration of validation loss, memory usage, and training time across various training methods for fine-tuning LLaMA-2 on the Alpaca dataset. BlockLLM demonstrates superior performance, achieving lower memory consumption and reduced training time.
@@
Existing strategies for memory-efficient training. To address these challenges, multiple strategies are being explored to reduce the number of parameters, gradients, and the corresponding optimizer state size. One popular strategy is the application of pruning methods, where a large set of parameters or entire layers are removed from the model architecture Wang et al. (2019); Ma et al. (2023); Sun et al. (2023). However, pruning approaches often require extensive retraining to recover lost accuracy Fan et al. (2019). Furthermore, identifying which parameters are crucial before training is challenging Michel et al. (2019); Sajjad et al. (2023). This challenge complicates implementation and can lead to generalization issues, particularly on diverse or unseen data Ma et al. (2023).
@@
PEFT (Parameter-Efficient-Fine-Tuning) methods Hu et al. (2021); Lialin et al. (2023); Hu et al. (2023) achieve memory efficiency by introducing low-rank matrices to the transformer architecture. This significantly reduces the number of trainable parameters needed during fine-tuning. Although integrating these low-rank matrices alleviates the extensive retraining demanded by pruning techniques, they can alter the training dynamics. This could potentially lead to quality issues during the merging phase He et al. (2021). The low-rank assumption may also constrain the model’s expressiveness, limiting its ability to fully capture complex patterns in the data. Furthermore, the additional parameters introduced by PEFT methods can increase the model’s parameter size, countering efforts to reduce overall model size.
@@
A recent work, GaLore Zhao et al. (2024) focuses on full parameter training and achieves memory efficiency by performing low-rank factorization of the gradients in specific layers. However, GaLore does not achieve high memory efficiency across all model types, as its gradient factorization method can only be applied to layers that satisfy the reversibility property. This limitation restricts its applicability and efficiency in models where not all layers exhibit this property.
@@
Techniques such as gradient and activation checkpointing Chen et al. (2016), quantization Han et al. (2015), and parameter offloading Rhu et al. (2016) are also commonly used to achieve memory savings. However, these methods often come with trade-offs such as increased computational overhead or compromised performance. For instance, checkpointing Chen et al. (2016) reduces memory usage but requires re-computation, quantization lowers precision and can affect accuracy Han et al. (2015) and parameter offloading increases data transfer latency Rhu et al. (2016). While these methods have their own limitations, many of these techniques are complementary to the approach presented in this work and provides additional opportunities for memory reduction when used in combination.
Block coordinate descent (BCD). BCD is popular algorithm in the large-scale optimization literature. At any training iteration $t$, instead of updating all the parameters $W$, BCD updates only a block of parameters $b_{t}$ by setting $W^{b_{t}}_{t+1}=W^{b_{t}}_{t}+d_{t}$, where $d_{t}$ is the update for that block. Importantly, $b_{t}$ does not stay fixed across iterations. As a result, BCD does not constrain model performance, which is often the case with low-rank approximation methods. Moreover, BCD doesn’t alter the model architecture in any way and preserves its structure throughout the training process. As it can be seen, this formulation directly falls in the reduced parameter training regime. This insight forms the cornerstone of our approach and hence the name BlockLLM.
A related study by (Belilovsky et al., 2019) demonstrated that sequentially solving one-hidden-layer problems could match the performance of large model training, inspiring us to tackle parameter and memory efficiency by training large models in parts. While their approach loses the full model context by focusing on smaller sub-problems, it inspired us to investigate how maintaining the full model context during training could potentially yield even better results. Finally, the convergence of BCD has been theoretically proven on various problem architectures in the optimization literature (Ramesh et al., 2023; Zeng et al., 2019; Nutini et al., 2022; Nesterov, 2012; Richtárik & Takáč, 2014; Shalev-Shwartz & Zhang, 2013), which inspired us to adapt it to large-scale LLM training.
Some previous works have explored training various neural network models using BCD (Zeng et al., 2019; Lau et al., 2018; Massart & Abrol, 2022). A recent parallel study (Luo et al., 2024) extends BCD to LLMs, focusing on memory efficient training. These methods update one block of parameters per iteration, chosen either randomly or cyclically. This approach is often inefficient for large-scale models because, in most cases only a small subset of parameters requires to be updated during finetuning (see our analysis 2 for details). If these critical parameters are not updated frequently enough, training becomes prolonged, impacting overall efficiency. In this work, we address this issue by updating parameters that are important for training more frequently. Our experimental results (see Figure 1) demonstrate that this strategy leads to both better performance and more efficient training. The offical code can be found here: https://github.com/RAmruthaVignesh/blockllm.
@@
The key contributions of this work are as follows:
  1. 1.
    We propose a novel parameter and memory-efficient algorithm, BlockLLM, where we dynamically select and train a block of parameters. This approach minimizes memory consumption by maintaining gradient and optimizer states only for the selected parameters.
  2. 2.
    We introduce a novel block selection criterion tailored for LLM training, where impactful parameters are updated more frequently. This leads to faster training, as important parameters are updated earlier in the process.
  3. 3.
    We demonstrate that BlockLLM achieves state-of-the-art training and generalization performance in both fine-tuning and pretraining tasks, while also enabling faster training and reduced memory usage.
  4. 4.
    We provide extensive ablation studies showing that BlockLLM is robust across various hyperparameter settings, including sparsity and patience.
@@ 2 Methodology
They key idea behind BlockLLM is to select and update only a subset of parameters during training, enabling us to achieve significant memory savings. An illustration of the method is given in Figure 2. However, the criteria for selecting the “right” subset of parameters is not clear. In this section, we look at magnitude pruning as a tool to identify parameter importance in the finetuning setting.
Figure 2: Given a large language model consisting of many layers, BlockLLM first finds the layers with the largest gradient $||\tilde{G}_{l}||$ (highlighted in orange) and selects a subset of the parameters. During optimization, only the selected layers will be updated ($\theta$) and the optimizer will keep track of its optimizer states only for those selected parameters (shown in orange).
@@
Magnitude pruning is a widely recognized technique for reducing the parameter count in neural network models Gupta et al. (2022). In this analysis, we use the weight magnitude of a parameter as a measure of parameter importance and study the impact of training on the selected parameters. First, we trained DistilBERT (Sanh et al., 2019) on the IMDb dataset (Maas et al., 2011) for sequence classification, achieving an accuracy of 92.02%. We then conducted inference on the GLUE-CoLA dataset (Wang et al., 2018) without fine-tuning, resulting in a significant drop in accuracy to $47.74\%$. This drop in performance, possibly due to domain shift, encouraged us to use this setup for our analysis.
Next, we performed magnitude pruning on the IMDb pretrained model at various sparsity levels. We then finetuned these pruned models on the GLUE-CoLA dataset. Let $s$ denote the sparsity level, $W^{t}$ represent the model parameters at iteration $t$, and $n$ be the total number of parameters. For each parameter $w_{i}$ where $i=1,\ldots,n$, we compute $|w_{i}|$. During training, we update only $S=\text{Top}_{k}|W^{0}|$, where $k=n\times(1-s)$. The results of these experiments, detailing the relationship between sparsity and accuracy, are summarized in Table 2.
@@
To further understand the effects of this parameter selection strategy, we analyzed $|W|$ before and after training. This helped understand which weights are updated more frequently and their impact on model performance. Specifically, we compared the initial weights $|W^{0}|$ and final weights $|W^{t}|$ of the model after fine-tuning on the CoLA dataset (Wang et al., 2018). Our findings are presented in Figure 3. The histogram on the left of Figure 3 shows $|w_{i}^{t}|$ for all $i$ where $\delta_{i}>\eta$. Here, $\delta_{i}=|w_{i}^{0}-w_{i}^{t}|$ and $\eta$ is the threshold. The histogram on the right of Figure 3 plots the frequency of $\delta$ in the updates, revealing that a large percentage of the updates were minor.
@@
(a) Histogram of the parameters that changed during finetuning.
(b) Histogram of changes in parameter magnitudes, revealing that most changes are small.
Figure 3: Analysis of weight magnitudes of DistilBERT model pretrained on IMDb dataset and finetuned on CoLA dataset suggesting that finetuning predominantly affects a narrow set of impactful parameters. This supports our hypothesis that focusing updates on a smaller number of important parameters can yield efficient training.
@@ 2.1 Analysis of Reduced Parameter Training
We conducted this experiment using the same setup: finetuning the DistilBERT model on the GLUE-CoLA dataset after pre-training on the IMDb dataset. The results of this experiment are summarized in the Table 3. Additionally, we performed similar experiments on other GLUE (Wang et al., 2018) datasets, with the results presented in Section A.3
Our observations indicate that as $s$ decreases, that is when more parameters are updated per iteration, the number of unique parameters $q$ increases. This results in better model performance. The increasing number of unique updates indicates that different parameters may become important at different stages of training, necessitating a more dynamic approach to parameter updates.
@@
Parameter Selection Criteria: Our analysis in Figure 3 indicates that the model frequently updates parameters with lower weight magnitudes. This contradicts the very premise of magnitude pruning. Therefore its not clear if adopting magnitude as a parameter importance criteria will help, in general. In our work, we bank on evidence from prior work on greedy parameter selection strategy Ramesh et al. (2023); Nutini et al. (2022) and use gradient as the parameter importance criteria.
Parameter Selection Frequency: As discussed earlier, determining when to change the set of parameters to update for a given training phase is crucial. In BlockLLM, we developed an adaptive strategy using the current loss value to decide when the parameter selection needs to be revisited.
2.2 BlockLLM
In this section we introduce BlockLLM, a parameter and memory efficient training method designed to reduce the number of trainable parameters in large language models (LLMs) without compromising training performance. Akin to other parameter-efficient fine-tuning (PEFT) methods such as LoRA (Hu et al., 2021) and ReLoRA (Lialin et al., 2023), BlockLLM updates only $k$ parameters at any iteration $t$. Here $k\ll n$ and $n$ is the total number of parameters. However, the main difference is that BlockLLM optimizes parameter selection by focusing on the most impactful parameters at different stages of the training process. The overall algorithm of BlockLLM is given in Algorithm 1.
@@ 2.2 BlockLLM
In the context of LLMs, at iteration $t$, the update to the parameter is the processed gradient $\tilde{G}_{t}$, calculated using optimizers such as Adam (Kingma & Ba, 2014). Specifically, for any layer $l$ of the model, the update is given by,
[…]
All the operations in equation 1 are applied element-wise. Here, $\beta_{1}$ and $\beta_{2}$ are hyperparameters of the optimizer. $G_{t}^{l}\in\mathbb{R}^{p\times q}$ is the gradient at layer $l$. $M_{t}$ and $V_{t}$ denote the bias-corrected first moment estimate and bias-corrected second moment estimate respectively.
BlockLLM achieves memory savings by storing these optimizer states $M_{t}$ and $V_{t}$ only for the currently selected layers, rather than for all parameters in the model. When the set of selected layers changes, the optimizer is reset with these new layers. This means it no longer maintains the optimizer states for the previously selected layers, similar to methods such as ReLoRA Lialin et al. (2023). As an alternative, we also tried to offload $M_{t}$ and $V_{t}$ of the selected parameters to CPU and re-load them as needed. But that did not improve the model performance. Thus we decided to adopt the former strategy to avoid the offloading operation in the interest of faster training.
@@
During the backward pass, BlockLLM selects the layers with large $||\tilde{G}_{t}^{l}||$ and updates only those layers. These selected layers are denoted as the set $S$. Note that by selecting full layers, we may not achieve the desired sparsity level $s$. Therefore, for each selected layer we construct a binary mask to retain only the top $k$ parameters by gradient magnitude:
$\text{mask}[i,j]=\begin{cases}1&\text{if }|\tilde{G}_{t}^{l}[i,j]|\geq\tau\\ 0&\text{otherwise},\end{cases}$
where $\tau$ is an estimated threshold, computed by looking at the gradient values of each layer. Specifically, $\tau$ is obtained computing the $(1-\zeta)^{th}$ percentile in $\tilde{G}_{t}^{l}$. The value of $\zeta$ is defined as $\zeta=(\Sigma_{p}-n_{s})/n_{s}$ (refer to Algorithm 2 for definitions of $\Sigma_{p}$ and $n_{s}$). Then, in every iteration $t$, the selected parameters in layers $l\in S$ are updated using the computed masks. The update rule is given by $W_{t+1}^{l}=W_{t}^{l}-\eta\left(\text{mask}\odot\tilde{G}_{t}^{l}\right)$ , where $\eta$ is the learning rate. An illustration of the proposed parameter selection procedure is given in Figure 2.
@@
However, there is one caveat with this approach. In the initial training iterations, the gradient estimates are known to be noisy. Additionally, in cases such as pretraining and finetuning with significant domain shifts, there is often very little useful inductive bias. Therefore, using gradients to select important parameters may prove to be detrimental to our cause in the initial few iterations.
To address this challenge, our experiments incorporate layer visit frequency $f$ into the selection criteria. Specifically, for any layer $l\in L=\{1,2,\ldots,n\}$, then $f_{l}$ represents the sum normalized number of times the layer has been selected. That is,
$S_{t}^{l}=\begin{cases}1&\text{if layer }l\text{ is selected at time }t\\ 0&\text{otherwise}.\end{cases}$
@@
The layer visit frequency $f_{l}$ for layer $l$ after $T$ time steps is given by $f_{l}=\frac{1}{T}\sum_{t=1}^{T}S_{t}^{l}$. Consequently, the layer selection criterion is modified to $|\tilde{G}_{l}|/f_{l}$. This modification favors layers with high gradient norms while also giving priority to layers that have been selected less frequently in previous iterations. Our experimental results demonstrate that this refined criterion enhances performance.
@@
The natural next question is how many iterations to update the parameters in the same set of layers $S$. BlockLLM addresses this by using the loss $\phi$ as a critical signal for determining when to change the parameter selection. Specifically, BlockLLM introduces a hyperparameter, patience $m$. At any iteration $t$, if $\phi_{t}$ equals to or exceeds the moving average of losses over the last $m$ iterations, the set $S$ is revised. The detailed parameter selection frequency algorithm is provided in Algorithm 2.
@@
The memory benefits of training with BlockLLM, stems primarily from its parameter-efficient training approach. In practice, updating fewer parameters directly reduces the number of gradients and optimizer states that need to be stored in VRAM. With $s\%$ sparsity, BlockLLM reduces the optimizer states by $s\%$ compared to full parameter training. Our empirical analysis corroborates this.
The parameter selection in BlockLLM relies on $\|\tilde{G}_{l}\|$ for each layer $l\in L$, which requires computing $\tilde{G}_{l}$ for all layers. This operation consumes significant memory required to store all the gradients. To reduce this, we sample a small number of $p$ additional layers per iteration besides the current block. Gradients for these $p$ layers are computed and their norms stored in a dictionary, which is then used for efficient parameter selection.
@@
10:         $H=[]$ //Reset loss history
@@
12:     // Update selected params
@@
15:         $M_{t}^{l},V_{t}^{l},\tilde{G}_{t}^{l}$ = Adam($M_{t-1}^{l},V_{t-1}^{l},G_{t-1}^{l}$)
16:         $\tilde{G}^{l}_{t}=\text{mask}\odot\tilde{G}_{t}^{l}$
17:         $W_{t+1}^{l}=W_{t}^{l}-\eta\tilde{G}_{t}^{l}$
18:     end for
@@
11:     for each $l\in S$ do
12:         Let $\text{mask}_{l}$ $=\mathbf{0}_{|G_{t}^{l}|}$
13:         for each $i,j$ in mask do
14:              if $\|\tilde{G}_{t}^{l}[i,j]\|\geq\tau$ then
15:                  $\text{mask}_{l}[i,j]=1$
16:              end if
17:         end for
@@
19:     return mask, $S$
20:end function